I have twice the age you had when I had the age you have. When you will be as old as I am, the sum of our ages will be 72 years. What are our ages ??


1
Expert's answer
2021-08-18T16:09:13-0400

let 's assume

I age =x

and your age =y

Gives

2x=y

x+y=72

Put value

x+2x=72

3x=72

x=24year

Put x value and find y value

24+y=49

y=49-24=48year

y=48year

I age x=24year